OpenStudy (anonymous):

let v be the space spanned by v1=cos^2x,v2=sin^2x,v3=sin2x.find a basis for v

OpenStudy (amistre64):

sadly ive never had to use trig functions as vectors :/

OpenStudy (amistre64):

sin2x = 2 sin(x) cos(x) sin^2 = 1/2 - cos(2x)/2 cos^2 = 1/2 + cos(2x)/2 .... yeah, nothing comes to mind from that